# Gradient Cuff: Detecting Jailbreak Attacks on Large Language Models by Exploring Refusal Loss Landscapes

### v2

@@ 4.1 Experiment Setup
Malicious User Queries. We sampled 100 harmful behavior instructions from AdvBench  [^2] https://github.com/llm-attacks/llm-attacks/blob/main/data/advbench/harmful_behaviors.csv in (Zou et al., 2023) as jailbreak templates, each to elicit the target LLM to generate certain harmful responses. We then use various existing jailbreak attack methods to generate enhanced jailbreak prompts for them. Specifically, for each harmful behavior instruction, we use GCG (Zou et al., 2023) to generate a universal adversarial suffix, use AutoDAN (Liu et al., 2023), PAIR (Chao et al., 2023), and TAP (Mehrotra et al., 2023) to generate a new instruction, use LRL (Yong et al., 2023) to translate it into low source languages that rarely appear in the training phase of the target LM such as German, Swedish, French and Chinese, and use Base64 (Wei et al., 2023a) to encode them in base64 format. See Appendix A.2 for more details on generating jailbreak prompts. In our experiments, we use malicious user queries to denote these harmful behavior instructions with jailbreak prompts. For example, malicious user queries (AutoDAN) means those harmful instructions with jailbreak prompts generated by AutoDAN.
@@
Aligned LLMs. We conduct the jailbreak experiments on 2 aligned LLMs: LLaMA-2-7B-Chat (Touvron et al., 2023) and Vicuna-7B-V1.5 (Zheng et al., 2023). LLaMA-2-7B-Chat is the aligned version of LLAMA-2-7B. Vicuna-7B-V1.5 is also based on LLAMA2-7B and has been further supervised fine-tuned on 70k user-assistant conversations collected from ShareGPT  [^4] https://sharegpt.com. We use protected LLM to represent these two models in the experiments.
Defense Baselines. Wecompare our method with various jailbreak defense methods including PPL (Jain et al., 2023), Erase-check (Kumar et al., 2023), SmoothLLM (Robey et al., 2023), and Self-Reminder (Robey et al., 2023). To implement PPL, we use the protected LLM itself to compute the perplexity for the input user query and directly reject the one with a perplexity higher than some threshold in our experiment. For Erase-Check, we employ the LLM itself to serve as a safety checker to check whether the input query or any of its erased sub-sentences is harmful. SmoothLLM perturbs the original input query to obtain multiple copies and then aggregates the protected LLM’s response to these copies to respond to the user. Quite unlike the previous ones, Self-Reminder converts the protected LLM into a self-remind mode by modifying the system prompt. For more details on the implementation of these baselines, please refer to Appendix A.7.
@@ 4.2 Performance Evaluation and Comparison with Existing Methods
We begin by evaluating our method as well as all the baselines except Self-Reminder against 6 different jailbreak attacks (GCG, AutoDAN, PAIR, TAP, Base64, and LRL) and benign user queries. We report the average refusal rate across these 6 malicious user query datasets as True Positive Rate (TPR) and the refusal rate on benign user queries as False Positive Rate (FPR). From Figure LABEL:fig:baseline_main we can summarize that Gradient Cuff stands out on both benign queries and malicious queries, attaining high TPR and low FPR. Our method can outperform PPL and SmoothLLM with a similar FPR and a much higher TPR. Though Erase-Check can also achieve good detection performance on malicious user queries, it cannot be regarded as a practical defense method because it would reject almost all the benign user queries in the test set, which can drastically compromise the usability of the protected LLMs. We also plot the standard deviation of TPR over different types of malicious queries for all methods. The results shown in Figure LABEL:fig:baseline_main_llama2 and LABEL:fig:baseline_main_vicuna demonstrate that our method has the most balanced performance across all types of jailbreaks considered in this paper. Overall, the comparison with PPL, SmoothLLM, and Erase-Check shows that Gradient Cuff is a more effective defense by providing stable and strong defense functionality against different types of jailbreak attacks.
@@
Language Model Defense Method FPR TPR LLaMA2-7B-Chat Self-Reminder $0.134$ $0.820$ Self-Reminder (GC) $0.126$ $0.920^{\blacklozenge}$ Gradient Cuff $0.070$ $0.968^{\bigstar}$ Vicuna-7B-V1.5 Self-Reminder $0.034$ $0.468$ Self-Reminder (GC) $0.044$ $0.646^{\blacklozenge}$ Gradient Cuff $0.026$ $0.669^{\bigstar}$
From the comparison, we can conclude that Gradient Cuff can significantly enhance Self-Reminder. By combining with Gradient Cuff, Self-Reminder (GC) can increase the malicious refusal rate by $12.20\%$ on LLaMA-2-7B-Chat and $\mathbf{38.03\%}$ on Vicuna-7B-V1.5. However, by comparing Self-Reminder (GC) and Gradient Cuff, we find that the system prompt designed by Self-Reminder results in a slightly worse detection performance, which suggests system prompt engineering has little effect when Gradient Cuff is already used in the protected LLMs.
@@ 4.3 Adaptive Attack
Adaptive attack is a commonly used evaluation scheme for defenses against adversarial attacks (Tramèr et al., 2020) with the assumption that the defense mechanisms are transparent to an attacker. Some studies on jailbreak defense also test their method against adaptive attacks (Robey et al., 2023; Xie et al., 2023). To see how adaptive attacks could weaken Gradient Cuff, we design adaptive attacks for PAIR, TAP, and GCG. Specifically, we design Adaptive-PAIR, Adaptive-TAP, and Adaptive-GCG to jailbreak protected LLMs equipped with Gradient Cuff. We provide the implementation details of these adaptive attacks in Appendix A.9. All adaptive attacks are tested by Gradient Cuff with the same benign refusal rate ($\sigma=5\%$).
As shown in Table 2, we find that Gradient Cuff is robust to Adaptive-GCG attack while the performance can be reduced by Adaptive PAIR and Adaptive TAP, especially when defending against Adaptive-PAIR on Vicuna-7B-V1.5, where the malicious refusal rate drops from $0.694$ to $0.356$.
@@
Table 2: Performance evaluation under adaptive attacks. The reported value is Gradient Cuff’s refusal rate against the corresponding jailbreak. The difference shows the change in performance when the adaptive attack is in place.
Language Model Jailbreak Adaptive Attack w/o w/ difference LLaMA-2-7B-Chat PAIR $0.770$ $0.778$ $+0.008$ TAP $0.950$ $0.898$ $-0.052$ GCG $0.988$ $0.986$ $-0.002$ Vicuna-7B-V1.5 PAIR $0.694$ $0.356$ $-0.338$ TAP $0.646$ $0.562$ $-0.084$ GCG $0.892$ $0.880$ $-0.012$
@@
We further compare our method with other defense baselines. Figure LABEL:fig:adaptive_attack_compare shows that our method is the best defense in terms of average restful rate on malicious queries. When protecting Vicuna-7B-V1.5, Gradient Cuff outruns SmoothLLM and PPL by $91.4\%$ and $81.6\%$ against Adaptive-PAIR while outperforming SmoothLLM and PPL by $52.7\%$ and $47.9\%$ against Adaptive-TAP. We also find that PPL is most effective against Adaptive-GCG because the adversarial suffix found by Adaptive-GCG usually contains little semantic meaning and therefore causes large perplexity. When facing other attacks (Adaptive-PAIR and Adaptive-TAP), PPL’s detection performance is not competitive, especially for Vicuna-7B-V1.5.
4.4 Query Budget Analysis
Recall that we have discussed the approximation error of the gradient estimation in Section 3.2. We conclude that we can decrease the approximation errors by choosing larger values for $N$ and $P$. However, in Section 3.3 we show that the total query time cost $C_{\theta}(N,P,x)=N\cdot(P+1)\cdot t_{\theta}(x)$ would also increase with $N$ and $P$. We did ablation studies on $N$ and $P$ in Appendix A.11 and found that it will have a better performance-efficiency trade-off when keeping $N$ fixed while changing $P$. Therefore, in this section, we fix $N=10$ and vary $P$ from $1$ to $10$ to evaluate Gradient Cuff under varying query times $q=N\cdot(P+1)$. As SmoothLLM also needs to query $T_{\theta}$ multiple times, we compare it with ours given the same query budget $q$.
The results in Figure LABEL:fig:ptimes show that when increasing the query budget, both SmoothLLM and our method would get better defense capability. The benign refusal rate for SmoothLLM is almost zero regardless of the number of query times, while the benign refusal rate for Gradient Cuff can be controlled by adjusting $\sigma$. Though SmoothLLM can maintain a similar benign refusal rate to Gradient Cuff when the $\sigma$ value is set to $5\%$, in terms of malicious refusal rate, Gradient Cuff outruns SmoothLLM by a large margin when the query times exceed $20$. For example, when allowing querying $T_{\theta}$ 110 times, Gradient Cuff can achieve 1.16×1.16$\times$1.16 × the malicious refusal rate of SmoothLLM on LLaMA2-7B-Chat and 1.92×1.92$\times$1.92 × the refusal rate on Vicuna-7B-V1.5.
@@ 4.5 Utility Analysis
In addition to the demonstrated improved defense capability, we further study how Gradient Cuff would affect the utility of the protected LLM. We compare the zero-shot performance of the Vicuna pair (Vicuna-7B-V1.5, Vicuna-7B-V1.5 with Gradient Cuff) and the LLaMA-2 pair (LLaMA-2-7b-Chat, LLaMA-2-7b-Chat with Gradient Cuff) on the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al., 2021).
Figure LABEL:fig:utility shows that Gradient Cuff does not affect the utility of the LLM on the non-rejected test samples. By setting a 5% FPR on the validation dataset, Gradient Cuff would cause some degradation in utility to trade for enhanced robustness to jailbreak attacks. We also note that the model developers can adjust the utility-robustness trade-off by controlling the detection threshold $t$ in Gradient Cuff.
@@ 5 Conclusion
In this paper, we define and study the refusal loss function of an LLM to exploit its discriminative functional properties for designing an effective jailbreak detection method called Gradient Cuff. Gradient Cuff features a two-step jailbreak detection procedure by sequentially checking the functional value and gradient norm of the refusal loss landscape. Our extensive experiments on 2 aligned LLMs (LLaMA-2-7b-Chat and Vicuna-7B-V1.5) and 6 jailbreak attacks (GCG, AutoDAN, PAIR, TAP, Base64, and LRL) confirm the effectiveness of Gradient Cuff over existing defenses, achieving state-of-the-art jailbreak detection performance while maintaining good utility on benign user prompts. One important future research direction is to further improve the detection performance against adaptive attacks and reduce the inference costs of Gradient Cuff.
Impact Statements
With the proliferation of LLMs and their applications across domains, reducing the risks of jailbreaking is essential for the creation of adversarially aligned LLMs. Our estimated impact is as broad as the LLMs themselves, since safety alignment is fundamental to the development and deployment of LLMs. We do not currently foresee any negative social impact from this work.

### v3

@@ 4.1 Experiment Setup
Malicious User Queries. We sampled 100 harmful behavior instructions from AdvBench [^2] https://github.com/llm-attacks/llm-attacks/blob/main/data/advbench/harmful_behaviors.csv in [33] as jailbreak templates, each to elicit the target LLM to generate certain harmful responses. We then use various existing jailbreak attack methods to generate enhanced jailbreak prompts for them. Specifically, for each harmful behavior instruction, we use GCG [33] to generate a universal adversarial suffix, use AutoDAN [17], PAIR [4], and TAP [18] to generate a new instruction, use LRL [30] to translate it into low source languages that rarely appear in the training phase of the target LM such as German, Swedish, French and Chinese, and use Base64 [26] to encode them in base64 format. See Appendix A.2 for more details on generating jailbreak prompts. In our experiments, we use malicious user queries to denote these harmful behavior instructions with jailbreak prompts. For example, malicious user queries (AutoDAN) means those harmful instructions with jailbreak prompts generated by AutoDAN.
@@
(a) LLaMA2-7B-Chat
(b) Vicuna-7B-V1.5
Figure 2: Performance evaluation on LLaMA2-7B-Chat (a) and Vicuna-7B-V1.5 (b). The horizon axis represents the refusal rate of benign user queries (FPR), and the vertical axis shows the average refusal rate across 6 malicious user query datasets (TPR). The error bar shows the standard deviation between the refusal rate of these 6 jailbreak datasets. We also report the MMLU accuracy of Low-FPR methods to show their utility. Complete results can be found in Appendix A.9.
Aligned LLMs. We conduct the jailbreak experiments on 2 aligned LLMs: LLaMA-2-7B-Chat [23] and Vicuna-7B-V1.5 [32]. LLaMA-2-7B-Chat is the aligned version of LLAMA-2-7B. Vicuna-7B-V1.5 is also based on LLAMA2-7B and has been further supervised fine-tuned on 70k user-assistant conversations collected from ShareGPT [^4] https://sharegpt.com. We use protected LLM to represent these two models in the experiments.
Defense Baselines. We compare our method with various jailbreak defense methods including PPL [12], Erase-check [14], SmoothLLM [21], and Self-Reminder [28]. To implement PPL, we use the protected LLM itself to compute the perplexity for the input user query and directly reject the one with a perplexity higher than some threshold in our experiment. For Erase-Check, we employ the LLM itself to serve as a safety checker to check whether the input query or any of its erased sub-sentences is harmful. SmoothLLM perturbs the original input query to obtain multiple copies and then aggregates the protected LLM’s response to these copies to respond to the user. Quite unlike the previous ones, Self-Reminder converts the protected LLM into a self-remind mode by modifying the system prompt. Though we mainly focus on unsupervised methods, we also conclude comparisons with two supervised methods: LLaMA-Guard [22] and Safe-Decoding [29]. We use the LLaMA-Guard-2-8B to implement LLaMA-Guard. For more details on the implementation of these baselines, please refer to Appendix A.8.
@@ 4.2 Performance Evaluation and Comparison with Existing Methods
We begin by evaluating our method as well as all the baselines except Self-Reminder against 6 different jailbreak attacks (GCG, AutoDAN, PAIR, TAP, Base64, and LRL) and benign user queries. We report the average refusal rate across these 6 malicious user query datasets as True Positive Rate (TPR) and the refusal rate on benign user queries as False Positive Rate (FPR). From Figure 2 we can summarize that Gradient Cuff stands out on both benign queries and malicious queries, attaining high TPR and low FPR. Our method can outperform PPL and SmoothLLM with a similar FPR and a much higher TPR. Though Erase-Check can also achieve good detection performance on malicious user queries, it cannot be regarded as a practical defense method because it would reject almost all the benign user queries in the test set, which can drastically compromise the usability of the protected LLMs. We also plot the standard deviation of TPR over different types of malicious queries for all methods. The results shown in Figure 2(a) and 2(b) demonstrate that our method has the most balanced performance across all types of jailbreaks considered in this paper.
Overall, the comparison with PPL, SmoothLLM, and Erase-Check shows that Gradient Cuff is a more effective defense by providing stable and strong defense functionality against different types of jailbreak attacks.
@@
Language Model Defense Method FPR TPR
LLaMA2-7B-Chat Self-Reminder $0.134\pm 0.034$ $0.820\pm 0.016$
Self-Reminder (GC) $0.126\pm 0.033$ $0.920\pm 0.011^{\blacklozenge}$
Gradient Cuff $0.070\pm 0.026$ $0.968\pm 0.007^{\bigstar}$
Vicuna-7B-V1.5 Self-Reminder $0.034\pm 0.018$ $0.472\pm 0.020$
Self-Reminder (GC) $0.044\pm 0.021$ $0.637\pm 0.020^{\blacklozenge}$
Gradient Cuff $0.026\pm 0.016$ $0.665\pm 0.019^{\bigstar}$
From the comparison, we can conclude that Gradient Cuff can significantly enhance Self-Reminder. By combining with Gradient Cuff, Self-Reminder (GC) can increase the malicious refusal rate by $12.20\%$ on LLaMA-2-7B-Chat and $\mathbf{34.96\%}$ on Vicuna-7B-V1.5. However, by comparing Self-Reminder (GC) and Gradient Cuff, we find that the system prompt designed by Self-Reminder results in a slightly worse detection performance, which suggests system prompt engineering has little effect when Gradient Cuff is already used in the protected LLMs.
For completeness, we compared with two supervised methods: LLaMA-Guard and Safe Decoding. Table 2 shows that LLaMA-Guard achieves comparable TPR results with Gradient Cuff on Vicuna-7B-V1.5 but a much lower TPR performance on LLaMA2-7B-Chat. Though Safe-Decoding achieves the largest average TPR against jailbreak attacks, its FPR is much higher, bringing notable utility degradation. Though LLaMA-Guard is a model-agnostic method, it shows model-specific results because we use different jailbreak prompts to evaluate it on different LLMs. The experimental results showed that the Gradient Cuff consistently stands out even when compared with supervised methods.
Table 2: Performance comparison with supervised methods.
Language Model Defense Method FPR TPR
LLaMA2-7B-Chat Gradient Cuff $0.022\pm 0.015$ $0.883\pm 0.013$
LLaMA-Guard $0.040\pm 0.020$ $0.760\pm 0.017$
Safe-Decoding $0.080\pm 0.027$ $0.955\pm 0.008$
Vicuna-7B-V1.5 Gradient Cuff $0.034\pm 0.018$ $0.743\pm 0.018$
LLaMA-Guard $0.040\pm 0.020$ $0.773\pm 0.017$
Safe-Decoding $0.280\pm 0.045$ $0.880\pm 0.013$
4.3 Effectiveness of Gradient Norm Rejection
We validate the effectiveness and necessity of using Gradient Norm Rejection as the second detection stage in Gradient Cuff by comparing the performance between Gradient Cuff and Gradient Cuff (w/o 2nd stage). Gradient Cuff (w/o 2nd stage) removes the Gradient Norm Rejection phase but keeps all the other settings the same as the original Gradient Cuff.
From Table 3, we can find that by adding the second stage and setting $\sigma$ to 1%, the TPR can be improved by a large margin (+0.099 on LLaMA2 and +0.240 on Vicuna), while the FPR is almost not changed (+0.000 on LLaMA2 and +0.002 on Vicuna). When we adjust the threshold in stage 2 by changing the $\sigma$ value from 1% to 5%, the performance gains in TPR can be further improved. These results verify the effectiveness of the Gradient Norm Rejection step in Graient Cuff.
Table 3: Performance evaluation of Gradient Cuff and Gradient Cuff (w/o 2nd stage). We remove the second stage or adjust the detection threshold of the 2nd stage to show its significance.
Language Model Defense Method FPR TPR
LLaMA2-7B-Chat Gradient Cuff (w/o 2nd stage) $0.012\pm 0.011$ $0.698\pm 0.019$
Gradient Cuff ($\sigma=1\%$) $0.012\pm 0.011$ $0.797\pm 0.016$
Gradient Cuff ($\sigma=5\%$) $0.022\pm 0.015$ $0.883\pm 0.013$
Vicuna-7B-V1.5 Gradient Cuff (w/o 2nd stage) $0.008\pm 0.009$ $0.296\pm 0.019$
Gradient Cuff ($\sigma=1\%$) $0.010\pm 0.010$ $0.536\pm 0.020$
Gradient Cuff ($\sigma=5\%$) $0.034\pm 0.018$ $0.743\pm 0.018$
@@ 4.4 Adaptive Attack
Adaptive attack is a commonly used evaluation scheme for defenses against adversarial attacks [24] with the assumption that the defense mechanisms are transparent to an attacker. Some studies on jailbreak defense also test their method against adaptive attacks [21, 28]. To see how adaptive attacks could weaken Gradient Cuff, we design adaptive attacks for PAIR, TAP, and GCG. Specifically, we design Adaptive-PAIR, Adaptive-TAP, and Adaptive-GCG to jailbreak protected LLMs equipped with Gradient Cuff. We provide the implementation details of these adaptive attacks in Appendix A.10. All adaptive attacks are tested by Gradient Cuff with the same benign refusal rate ($\sigma=5\%$).
@@
Table 4: Performance evaluation under adaptive attacks. The reported value is Gradient Cuff’s refusal rate against the corresponding jailbreak attack.
Language Model Jailbreak Adaptive Attack
w/o w/
LLaMA-2-7B-Chat PAIR $0.770\pm 0.042$ $0.778\pm 0.042$
TAP $0.950\pm 0.022$ $0.898\pm 0.030$
GCG $0.988\pm 0.011$ $0.986\pm 0.012$
Vicuna-7B-V1.5 PAIR $0.694\pm 0.046$ $0.356\pm 0.048$
TAP $0.570\pm 0.050$ $0.562\pm 0.050$
GCG $0.892\pm 0.031$ $0.880\pm 0.032$
Table 5: Robustness-Utility evaluation on MMLU benchmark.
Language Model Methods TPR MMLU Accuracy
LLaMA-2-7B-Chat Gradient Cuff $0.883\pm 0.013$ $0.378$
SmoothLLM $0.763\pm 0.017$ 0.271
PPL $0.732\pm 0.018$ 0.411
Vicuna-7B-V1.5 Gradient Cuff $0.743\pm 0.018$ 0.462
SmoothLLM $0.413\pm 0.020$ 0.303
PPL $0.401\pm 0.020$ 0.487
As shown in Table 4, Gradient Cuff is robust to Adaptive-GCG attack while the performance can be mildly reduced by Adaptive PAIR and Adaptive TAP, especially when defending against Adaptive-PAIR on Vicuna-7B-V1.5, where the malicious refusal rate drops from $0.694$ to $0.356$.
@@
We further compare our method with other defense baselines. Figure 3 shows that our method is the best defense in terms of the average refusal rate on malicious queries. On Vicuna-7B-V1.5, Gradient Cuff outruns SmoothLLM and PPL by $91.4\%$ and $81.6\%$ against Adaptive-PAIR while outperforming SmoothLLM and PPL by $52.7\%$ and $47.9\%$ against Adaptive-TAP. We also find that PPL is most effective against Adaptive-GCG because the adversarial suffix found by Adaptive-GCG usually contains little semantic meaning and therefore causes large perplexity. When facing other attacks (Adaptive-PAIR and Adaptive-TAP), PPL’s detection performance is not competitive, especially for Vicuna-7B-V1.5. In Appendix A.11, we validated the effect of adaptive attacks against Gradient Cuff by showing that they intended to decrease the norm of the refusal loss gradient.
(a) LLaMA2-7B-Chat
(b) Vicuna-7B-V1.5
Figure 3: Performance comparison against adaptive jailbreak attacks.
@@ 4.5 Utility Analysis
In addition to the demonstrated improved defense capability, we further study how Gradient Cuff would affect the utility of the protected LLM. We compare the zero-shot performance of the Vicuna pair (Vicuna-7B-V1.5 & Vicuna-7B-V1.5 with Gradient Cuff) and the LLaMA-2 pair (LLaMA-2-7b-Chat & LLaMA-2-7b-Chat with Gradient Cuff) on the Massive Multitask Language Understanding (MMLU) benchmark [9]. Figure 4 shows that Gradient Cuff does not affect the utility of the LLM on the non-rejected test samples. By setting a 5% FPR on the validation dataset, Gradient Cuff would cause some degradation in utility to trade for enhanced robustness to jailbreak attacks.
We also report the utility for existing baselines in Table 5. We find that though SmoothLLM can achieve a very low FPR as shown in Figure 2, it causes a dramatically large utility degradation because it has modified the user query, inevitably compromising the semantics of the query. PPL attains the best utility and Gradient Cuff achieves the best performance-utility trade-off by (a) keeping the comparable utility with PPL (b) attaining a much higher TPR than the best baselines (e.g., 0.743 vs 0.413 on Vicuna-7B-V1.5) against jailbreak attacks.
(a) LLaMA2-7B-Chat
(b) Vicuna-7B-V1.5
Figure 4: Utility evaluation on MMLU [9] (zero-shot) with and without Gradient Cuff.
5 Discussion
Extra Inference Cost. Experimental results in Appendix A.16 show that our method achieves the most effective trade-off between performance and inference efficiency when doing jailbreak defense. We also think this trade-off is inevitable (though it could be improved) yet acceptable, as users may be less incentivized to use a model/service if it does not have proper safety guardrails.
Jailbroken Assessment. Existing studies often rely on checking whether an LLM’s response contains certain predefined keywords or phrases to assess the jailbroken. However, this method has obvious limitations, as it is difficult to create an exhaustive list of phrases that could cover all possible jailbreaking scenarios. Consequently, we need a more reliable method to accurately identify successful jailbreaking attempts. In Appendix A.6, we use GPT-4 and LLaMA-2-Guard-8B to compute the AS. The results are consistent with the keyword-based ASR evaluations.
Application on close-sourced LLMs. When implementing our detection method, we adopt the white-box settings assuming the model weights and internal representations are available to the defender. Gradient Cuff is applicable to close-sourced LLMs if it is deployed by the model developer who has full access to the model’s parameters including the embedding layers.
@@ 6 Conclusion
In this paper, we define and study the refusal loss function of an LLM to exploit its discriminative functional properties for designing an effective jailbreak detection method called Gradient Cuff. Gradient Cuff features a two-step jailbreak detection procedure that sequentially checks the refusal loss landscape’s functional value and gradient norm. Our experiments on 2 aligned LLMs (LLaMA-2-7b-Chat and Vicuna-7B-V1.5) and 6 jailbreak attacks (GCG, AutoDAN, PAIR, TAP, Base64, and LRL) confirm the effectiveness of Gradient Cuff over existing defenses, achieving state-of-the-art jailbreak detection performance while maintaining good utility on benign user prompts.
Acknowledgments and Disclosure of Funding
This work was supported by the JC STEM Lab of Intelligent Design Automation funded by The Hong Kong Jockey Club Charities Trust for Xiaomeng Hu and Tsung-Yi Ho.